Cassandra has six different colored bracelets. In how many ways can she wear three bracelets at a time?
Answer by stanbon(75887)

Cassandra has six different colored bracelets. In how many ways can she wear three bracelets at a time?
-----------
# of sets of three: 6C3 = 6*5*4/1*2*3 = 20
--------------
orders in which the three can be worn = 3*2*1 = 6
-----------------
Cheers,
Stan H.